# Tree balance in phylogenetic models

### v2

@@ Abstract
Tree shape statistics, particularly measures of tree (im)balance, play an important role in the analysis of the shape of phylogenetic trees. With applications ranging from testing evolutionary models to studying the impact of fertility inheritance and selection, or tumor development and language evolution, the assessment of tree balance is crucial. Currently, a multitude of at least 30 (im)balance indices can be found in the literature, alongside numerous other tree shape statistics.
This diversity prompts essential questions: How can we minimize the selection of indices to mitigate the challenges of multiple testing? Is there a preeminent balance index tailored to specific tasks? Previous studies comparing the statistical power of indices in detecting trees deviating from the Yule model have been limited in scope, utilizing only a subset of indices and alternative tree models.
This research expands upon the examination of index power, encompassing all established indices and a broader array of alternative models. Our investigation reveals distinct groups of balance indices better suited for different tree models, suggesting that decisions on balance index selection can be enhanced with prior knowledge. Furthermore, we present the R software package poweRbal which allows the inclusion of new indices and models, thus facilitating future research.
@@ 1 Introduction
A key aspect in the study of evolution is to understand the forces that drive speciation and extinction processes and their effect on the evolution of species and taxa of higher level [16]. In the 1970s, the idea formed that such evolutionary processes can be modeled using stochastic processes (see, e.g., [39, 38, 45, 17, 37, 44]). Around the same time, researchers started assessing the shapes of (phylogenetic) trees via tree shape statistics (mostly tree balance indices, which measure the degree of imbalance/asymmetry in trees) as well as analyzing the probabilities of certain shape patterns under basic phylogenetic models [12, 18], and several biogeographic studies introduced the use of null models in macroevolution [41, 48, 47]. At the intersection of these fields, a whole research branch emerged, which is concerned with developing stochastic models that are good and useful explanations of (species) evolution based on the analysis of tree shapes (see, e.g., [43, 49, 29, 33, 34, 5, 7, 8, 52, 36, 41, 15, 1]). Intuitively, when the moments of a tree shape statistic under a given null model are known, or when the null distribution can be approximated through simulation, statistical methods to test a given tree against the null hypothesis (i.e., that the tree was obtained under the null model) can be devised. These results can then be used to decide whether a given null model is indeed a realistic model for evolution.
@@
The main purpose of our work, though, is to bridge these two pillars by providing the means to examine the power of all established tree shape statistics for a wide array of models. While we focus on their power to detect the Yule model against certain discrete-time models for the purpose of this paper, our general framework can easily be extended to other null and alternative models as well as new tree shape statistics. In particular, we introduce the R software package poweRbal, which can be used to perform analyses analogous to the ones presented here. This package currently includes all established tree balance indices as well as numerous macroevolutionary models, but is implemented in such a way that new tree shape statistics and new models can easily be incorporated. Detailed instructions on how to use the package are given in Section 5. In general, we suggest to use this software package as a precursor to subsequent analyses and research: The user specifies the research question (e.g., which null and alternative models to investigate), and inputs these choices into our software. Our software then provides the user with the most powerful tree shape statistics for this setting to be used in subsequent statistical analyses of the data. This minimizes problems related to multiple testing and ensures that no better (i.e., more powerful) tree shape statistic is overlooked.
@@
As indicated above, our paper is not the first to analyze tree shape statistics in the context of evolutionary models, but previous studies have been limited in scope. Nonetheless, we briefly summarize these studies and highlight some alternative approaches that have been used in the literature. First, similar to our study, several authors have studied the power of tree shape statistics to detect deviations from a given phylogenetic null model (e.g., [20, 29, 21, 1, 7, 22, 25, 26]). Using the Yule model as the null model, these studies have employed a handful of tree shape statistics and a handful of alternative models in their investigation. While some tree shape statistics such as the Colless index [10] have been analyzed in several of these studies, others, such as the rooted quartet index [11], have not been included at all. However, we remark that very recently and independent of our study, Khurana et al. [28] used a large selection of the tree shape statistics categorized in [13] in a study aimed at discriminating between empirical trees and trees generated under a constant-rate birth-death model. Nevertheless, by including all established tree shape statistics as well as a large variety of alternative models in our analyses, we complement and expand the existing literature.
Next to analyzing the power of tree shape statistics in the context of macroevolutionary null models, tree shape statistics have also been used in testing for biogeographic null models (see, e.g., [22, p. 108]). Here, the aim is to decide whether a local subphylogeny is significantly more symmetric or asymmetric than expected if a random phylogeny of the same size is picked from the underlying larger tree.
Finally, there have been some approaches of evaluating tree shape statistics outside the realm of classic statistical hypotheses tests. For instance, Matsen [31] took a geometric approach to quantifying the power of tree shape statistics to differentiate between similar and different trees, where similarity of trees was measured using the so-called nearest neighbor interchange distance. Intuitively, the idea is that under good tree shape statistics, similar trees should receive similar values, whereas distant trees should receive different values. Hayati et al. [19] recently followed up on Matsen’s idea, albeit using a different resolution function (for a more detailed discussion of both approaches, see [19, pp. 5-10]).
@@ 2 Preliminaries
Remark.
The leaves of a tree (sometimes also called “tips”) represent species or lineages (depending on the context), so we sometimes simplify explanations by saying, e.g., that “a leaf speciates” or “a leaf goes extinct”, which actually means that the species or lineage represented by that particular leaf undergoes said evolutionary event.
@@
A very intuitive example of how to measure tree imbalance is given by the Colless index, which compares the number of leaves in the two maximal pending subtrees (which are formally defined in the supplementary material [27]) of each inner node and takes the sum over these absolute differences (see Figure 1 on the right).
@@
In Tables 1 and 2 in the supplementary material [27] we list all tree (im)balance indices considered in this paper with a short description and their formal definition. We refer the reader to the respective references and [13] for further details.
@@
A very foundational tree model is the Yule model [54], which is a pure birth model, where all species have the same constant rate of speciation at all times. To obtain a tree shape under the Yule model using a forward process, starting with a single leaf, a leaf is chosen uniformly at random and replaced by a cherry (see Figure 1 on the left). Finally, a phylogenetic tree under the Yule model is obtained by assigning leaf labels uniformly at random to a shape generated by this process. The Yule model plays a central role in phylogenetics and has been used for many decades. It is known under a multitude of names (Yule model, equal-rates-Markov model (ERM), random branching model, Markovian dichotomous branching model, Yule-Harding model, Yule-Harding-Kingman model (YHK), or simply Markovian model) of which we will only use Yule model in the remainder of this manuscript. While it can be described under various approaches, the probabilities to obtain a specific (phylogenetic) tree can be explicitly stated as follows (see, e.g., [51, Proposition 1] or [50, Proposition 3.2]):
$P_{Y,n}(\mathcal{T})=\frac{2^{n-1}}{n!}\cdot\prod\limits_{v\in\mathring{V}(T)}\frac{1}{n_{v}-1}\qquad\text{and}\qquad P_{Y,n}(T)=2^{n-1-s(T)}\cdot\prod\limits_{v\in\mathring{V}(T)}\frac{1}{n_{v}-1}.$
While the Yule model is one of the oldest and one of the most famous tree models, and it is often used as a null model in phylogenetic analyses, there are many more models available to describe different evolutionary settings. Section 3 in the supplementary material [27] provides an extensive collection of tree models discussed in the literature.
Figure 1: Left: Illustration of the Yule model using the forward process (“tree growing”). At each step from $n$ to $n+1$, a leaf is chosen uniformly at random and replaced by a cherry. Thus, under the Yule model, the caterpillar tree on four leaves, occurs with probability $P_{4}(T^{cat}_{4})=\frac{2}{3}$, whereas the fully balanced tree of height two occurs with probability $P_{4}(T^{fb}_{2})=\frac{1}{3}$.

### v3

@@ Abstract
Tree shape statistics, particularly measures of tree (im)balance, play an important role in the analysis of the shape of phylogenetic trees. With applications ranging from testing evolutionary models to studying the impact of fertility inheritance and selection, or tumor development and language evolution, the assessment of measures of tree balance is important. Currently, a multitude of at least 30 (im)balance indices can be found in the literature, alongside numerous other tree shape statistics.
This diversity prompts essential questions: How can we assist researchers to choose only a small number of indices to mitigate the challenges of multiple testing? Is there a preeminent balance index tailored to specific tasks?
This research expands previous studies on the examination of index power, encompassing almost all established indices and a broader array of alternative models, such as a variety of trait-based models. Our investigation reveals distinct groups of balance indices better suited for different tree models, suggesting that decisions on balance index selection can be enhanced with prior knowledge. Furthermore, we present the R software package poweRbal which allows the inclusion of new indices and models, thus facilitating future research on the power of tree shape statistics.
@@ 1 Introduction
History of tree balance in phylogenetic models
A key aspect in the study of evolution is to understand the forces that drive speciation and extinction processes and their effect on the evolution of species and taxa of higher level [1]. In the 1960s and 1970s, the idea formed that such evolutionary processes can be modeled using stochastic processes (see, e.g., [2, 3, 4, 5, 6, 7, 8, 9]). Around the same time, researchers started assessing the shapes of (phylogenetic) trees via tree shape statistics (mostly tree balance indices, which measure the degree of imbalance/asymmetry in trees). Tree shape statistics are nowadays used in a variety of contexts, ranging from phylodynamics and the study of viruses (e.g., [10, 11, 12]) to phylolinguistics and the study of language diversification (e.g., [13]); see also [14, Chapter 27] for further applications of tree shape statistics. Returning to the context of macroevolution and the 1970s, the research community also started to analyze the probabilities of certain shape patterns under basic phylogenetic models [15, 16], and several biogeographic studies introduced the use of null models in macroevolution [17, 18, 19]. At the intersection of these topics, a whole research branch emerged, which is concerned with developing stochastic models that are good and useful explanations of diversification based on the analysis of tree shapes (see, e.g., [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 17, 30, 31]).
Intuitively, when the moments of a tree shape statistic under a given null model are known, or when the null distribution can be approximated through simulation – or sometimes through asymptotic distributions, statistical methods to test a given tree against the null hypothesis (i.e., that the tree was obtained under the null model) can be devised. These results can then be used to decide whether a given null model is indeed a realistic model for evolution.
@@
Aims of this manuscript
The main purpose of our work, though, is to bridge the above mentioned two pillars by providing the means to examine the power of all established tree shape statistics for a wide array of models. While we focus on their power to detect the Yule model against certain discrete-time models for the purpose of this paper, our general framework can easily be extended to other null and alternative models as well as new tree shape statistics. In particular, we introduce the R software package poweRbal, which can be used to perform analyses analogous to the ones presented here. This package currently includes all established tree balance indices as well as numerous macroevolutionary models, but is implemented in such a way that new tree shape statistics and new models can easily be incorporated. Detailed instructions on how to use the package are given in the supplementary material [40]. In general, we suggest to use this software package as a precursor to subsequent analyses and research: The user specifies the research question (e.g., which null and alternative models to investigate), and inputs these choices into our software. Our software then provides the user with the most powerful tree shape statistics for this setting to be used in subsequent statistical analyses of the data. This minimizes problems related to multiple testing and ensures that no better (i.e., more powerful) tree shape statistic is overlooked.
@@
As indicated above, our paper is not the first to analyze tree shape statistics in the context of evolutionary models, but previous studies have been limited in scope. Nonetheless, we briefly summarize these studies and highlight some alternative approaches that have been used in the literature. First, similar to our study, several authors have studied the power of tree shape statistics to detect deviations from a given phylogenetic null model (e.g., [32, 22, 33, 31, 26, 34, 35, 36]). Using the Yule model as the null model, these studies have employed a handful of tree shape statistics and a handful of alternative models in their investigation. While some tree shape statistics such as the Colless index [38] have been analyzed in several of these studies, others, such as the rooted quartet index [39], have not been included at all. However, we remark that very recently and independent of our study, Khurana et al. [42] used a large selection of the tree shape statistics categorized in [14] in a study aimed at discriminating between empirical trees and trees generated under a constant-rate birth-death model (see Section 6 in the supplementary material for a more thorough discussion of the study by Khurana et al. [42].) Nevertheless, by including all established tree shape statistics as well as a large variety of alternative models in our analyses, we complement and expand the existing literature.
Next to analyzing the power of tree shape statistics in the context of macroevolutionary null models, tree shape statistics have also been used in testing for biogeographic null models (see, e.g., [34, p. 108]). Here, the aim is to decide whether a local subphylogeny is significantly more symmetric or asymmetric than expected if a random phylogeny of the same size is picked from the underlying larger tree.
Finally, there have been some approaches of evaluating tree shape statistics outside the realm of classic statistical hypotheses tests. For instance, Matsen [43] took a geometric approach to quantifying the power of tree shape statistics to differentiate between similar and different trees, where similarity of trees was measured using the so-called nearest neighbor interchange distance. Intuitively, the idea is that under good tree shape statistics, similar trees should receive similar values, whereas distant trees should receive different values. Hayati et al. [44] recently followed up on Matsen’s idea, albeit using a slightly different approach (for a more detailed discussion of both approaches, see [44, pp. 5-10]).
@@ 2 Preliminaries
A rooted binary phylogenetic $X$-tree $\mathcal{T}$ is a tuple $\mathcal{T}=(T,\phi)$, where $T=(V(T),E(T))$ is a rooted tree with root $\rho$ and leaf set $V_{L}(T)$ such that each inner vertex $v\in V(T)\setminus V_{L}(T)$ has out-degree 2, and $\phi$ is a bijection from the leaf set $V_{L}(T)$ to $X$. Intuitively, $\mathcal{T}$ is obtained from $T$ by assigning labels or species names to the leaves of $T$. In particular, if $|X|=n$, we also have $|V_{L}(T)|=n$. $T$ is called a rooted binary tree and is also often referred to as the topology or tree shape of $\mathcal{T}$ [^1] Note that sometimes in the literature the term tree shape is used to refer to the distribution of edge lengths among internal and leaf edges, without specific reference to the topology [45]. However, we use the two terms topology and tree shape interchangeably as tree shape statistics is an established term for functions that assess the topology of trees. We use $\mathring{V}(T)$ to denote the set of inner vertices of $T$. For every $n\in\mathbb{N}_{\geq 1}$, we denote by $\mathcal{BT}^{\ast}_{n}$ the set of (isomorphism classes of) all rooted binary trees with $n$ leaves and by $\mathcal{BT}_{n}$ the set of (isomorphism classes of) all rooted binary phylogenetic $X$-trees with $n$ leaves.
Given a rooted binary tree $T$ and an arbitrary vertex $v\in V(T)$, we denote by $T_{v}$ the pending subtree of $T$ rooted in $v$ and use $n_{v}$ to denote the number of leaves in $T_{v}$. Let $v\in\mathring{V}(T)$ be an inner vertex of $T$ with children $v_{1}$ and $v_{2}$. We call $v$ a symmetry node or symmetry vertex if $T_{v_{1}}$ and $T_{v_{2}}$ are isomorphic, i.e., have the same shape. The number of symmetry nodes of $T$ is denoted as $s(T)$.
@@
In Tables 1 and 2 we list all tree (im)balance indices considered in this manuscript with a short description and their formal definitions. We refer the reader to the supplementary material [40] for further details on terminology.
A very intuitive example of how to measure tree imbalance is given by the Colless index (Table 2), which for each inner node $v$ compares the number of leaves in the two pending subtrees induced by $v$, say $n_{v_{1}}$ and $n_{v_{2}}$, and takes the sum over these absolute differences, also known as balance values $bal_{T}(v)=|n_{v_{1}}-n_{v_{2}}|$ (see Figure 1 on the right).
Table 1: List of tree balance indices (adapted from [14]). The first column contains the name of the index and a brief description. The second column contains its formal definition. Details on terminology can be found in the supplementary material [40].
Balance indices
Index Definition
$\boldsymbol{B_{1}}$ index [46] sum of the reciprocals of the heights of the subtrees of $T$ rooted at inner vertices of $T$ (except for $\rho$) $B_{1}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)\setminus\{\rho\}}h(T_{v})^{-1}$
$\boldsymbol{B_{2}}$ index [31, 44, 22, 46] Shannon-Wiener information function (measures the equitability of arriving at the leaves of $T$ when starting at the root and assuming equiprobable branching at each inner vertex) $B_{2}(T)\coloneqq-\sum\limits_{x\in V_{L}(T)}p_{x}\cdot\log(p_{x})$
Furnas rank [47, 22] rank of $T$ according to Furnas’ left-light rooted ordering on $\mathcal{BT}^{\ast}_{n}$ $F(T)\coloneqq r_{n}(T)$
Maximum width [11] maximum width at any depth $mW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)}w_{T}(i)$
Maximum width over maximum depth [11] ratio of maximum width and maximum depth $W/D(T)$ $\coloneqq\frac{mW(T)}{mD(T)}$
Modified maximum difference in widths [14, 11] maximum difference in widths of two consecutive depths $mdelW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)-1}{w_{T}(i+1)-w_{T}(i)}$
Rooted quartet index for binary trees [39] balance index based on the symmetry of the displayed quartets of $T$ $rQI(T)\coloneqq|\{Q\in\mathcal{Q}(T):Q\text{ has shape }T_{2}^{fb}\}|$
stairs2 [10, 11] ratio between the leaf numbers of the smaller and larger pending subtree over all inner vertices $st2(T)\coloneqq\frac{1}{n-1}\cdot\sum\limits_{v\in\mathring{V}(T)}\frac{\min\{n_{v_{1}},n_{v_{2}}\}}{\max\{n_{v_{1}},n_{v_{2}}\}}$
Table 2: List of tree imbalance indices (adapted from [14]). The first column contains the name of the index and a brief description. The second column contains its formal definition. Details on terminology can be found in the supplementary material [40].
Imbalance indices
Index Definition
Average leaf depth [37, 46] mean depth of the leaves of $T$ $\overline{N}(T)\coloneqq\frac{1}{n}\cdot\sum\limits_{x\in V_{L}(T)}\delta_{T}(x)$
Average vertex depth [48]; see also [49, 27, 50] mean depth of the vertices of $T$ $AVD(T)\coloneqq\frac{1}{|V|}\cdot\sum\limits_{v\in V(T)}\delta_{T}(v)$
Colijn-Plazzotta rank [51, 52] recursively defined bijective mapping between rooted binary trees and the positive integers $CP(T)\coloneqq 1$ for $T\in\mathcal{BT}_{1}^{\ast}$ and for $T=(T_{1},T_{2})\in\mathcal{BT}^{\ast}_{n}$ with $n\geq 2$, $CP(T)\coloneqq\frac{1}{2}CP(T_{1})(CP(T_{1})-1)+CP(T_{2})+1$ (with $CP(T_{1})\geq CP(T_{2})$)
Colless index [38] sum of the balance values of the inner vertices of $T$ $C(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{T}(v)=\sum\limits_{v\in\mathring{V}(T)}|n_{v_{1}}-n_{v_{2}}|$
Corrected Colless index [32] $C(T)$ divided by its maximum possible value $I_{C}(T)\coloneqq\frac{2\cdot C(T)}{(n-1)(n-2)}$
Equal weights Colless index [24] modification of $C(T)$ that weighs the imbalance values of all inner vertices of $T$ equally $I_{2}(T)\coloneqq\frac{1}{n-2}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}>2}\frac{bal_{T}(v)}{n_{v}-2}$
Total $\boldsymbol{I}$ and $\boldsymbol{I^{\prime}}$ index [30, 29] sum/total of the $I_{v}$ ($I_{v}^{\prime}$) values over all inner vertices $v$ with $n_{v}\geq 4$ $\Sigma I(T)\coloneqq\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I_{v}$ $\Sigma I^{\prime}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I^{\prime}_{v}$
Maximum depth [11] maximum depth of any vertex of $T$; equivalent to $h(T)$ $mD(T)\coloneqq\max\limits_{v\in V(T)}\delta_{T}(v)=h(T)$
Mean $\boldsymbol{I}$ and $\boldsymbol{I^{\prime}}$ index [30, 29] mean of the $I_{v}$ ($I_{v}^{\prime}$) values over all inner vertices $v$ with $n_{v}\geq 4$ $\overline{I}(T)\coloneqq\frac{1}{|\{v\in\mathring{V}(T):n_{v}\geq 4\}|}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I_{v}$ $\overline{I^{\prime}}(T)\coloneqq\frac{1}{|\{v\in\mathring{V}(T):n_{v}\geq 4\}|}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I^{\prime}_{v}$
Quadratic Colless index [53] sum of the squared balance values of the inner vertices of $T$ $QC(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{T}(v)^{2}$
Rogers $\boldsymbol{J}$ index [54] number of inner vertices of $T$ which are not perfectly balanced $J(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}(1-\mathcal{I}(bal_{T}(v)=0))$
$\boldsymbol{\widehat{s}}$-shape statistic [27] sum of $\log(n_{v}-1)$ over all inner vertices of $T$ $\widehat{s}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}\log(n_{v}-1)$
Sackin index [37, 46] sum of the leaf depths of $T$ $S(T)\coloneqq\sum\limits_{x\in V_{L}(T)}\delta_{T}(x)$
stairs1 [10, 11] proportion of inner vertices that are not perfectly balanced; normalization of Rogers $J$ index $st1(T)\coloneqq\frac{1}{n-1}\cdot\sum\limits_{v\in\mathring{V}(T)}(1-\mathcal{I}(bal_{T}(v)=0))$
Symmetry nodes index [36] number of inner vertices of $T$ that are not symmetry nodes $SNI(T)\coloneqq(n-1)-s(T)$
Total cophenetic index [55] sum of the cophenetic values of all distinct pairs of leaves of $T$ $\Phi(T)\coloneqq\sum\limits_{\begin{subarray}{c}\{x,y\}\in V_{L}(T)^{2}\\ x\neq y\end{subarray}}\varphi_{T}(x,y)$
Total internal path length [56] total of the depths of all inner vertices of $T$ $TIP(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}\delta_{T}(v)$
Total path length [57, 58, 59] total of the depths of all vertices of $T$ $TPL(T)\coloneqq\sum\limits_{v\in V(T)}\delta_{T}(v)$
Variance of leaf depths [60, 37, 46] variance of the depths of the leaves of $T$ $\sigma_{N}^{2}(T)\coloneqq\frac{1}{n}\cdot\sum\limits_{x\in V_{L}(T)}\left(\delta_{T}(x)-\overline{N}(T)\right)^{2}$
@@
A very foundational tree model is the Yule model [61], which is a pure birth model, where all species have the same constant rate of speciation at all times. To obtain a tree shape under the Yule model using a forward process, starting with a single leaf, a leaf is chosen uniformly at random and replaced by a cherry (see Figure 1 on the left). Finally, a phylogenetic tree under the Yule model is obtained by assigning leaf labels uniformly at random to a shape generated by this process. In particular, this means that for the generated tree shape $T\in\mathcal{BT}^{\ast}_{n}$, each of the possible $\frac{n!}{2^{s(T)}}$ phylogenetic trees is equally likely to be chosen. In practice, this can for instance be achieved by choosing, for $i=1\ldots,n$, one leaf at a time uniformly at random from the set of leaves that have not been labeled yet and assigning it label $i$ [^2] Note that this procedure leads to a multiset of phylogenetic trees in which precisely $2^{s(T)}$ many (isomorphic) versions of each phylogenetic tree are contained (due to the symmetry nodes). Thus, when drawing one phylogenetic tree uniformly at random from this multiset, every phylogenetic tree has the same probability. The Yule model plays a central role in phylogenetics and has been used for many decades. It is known under a multitude of names (Yule model, equal-rates-Markov model (ERM), random branching model, Markovian dichotomous branching model, Yule-Harding model, Yule-Harding-Kingman model (YHK), or simply Markovian model) of which we will only use Yule model in the remainder of this manuscript. While it can be described under various approaches, the probabilities to obtain a specific (phylogenetic) tree can be explicitly stated as follows (see, e.g., [62, Proposition 1] or [63, Proposition 3.2]):
$P_{Y,n}(\mathcal{T})=\frac{2^{n-1}}{n!}\cdot\prod\limits_{v\in\mathring{V}(T)}\frac{1}{n_{v}-1}\qquad\text{and}\qquad P_{Y,n}(T)=2^{n-1-s(T)}\cdot\prod\limits_{v\in\mathring{V}(T)}\frac{1}{n_{v}-1}.$ (1)
In particular the formula for $P_{Y,n}(\mathcal{T})$ has been extensively studied in the literature and connections to other fields have been established (see, e.g., [64, 65, 66]).
Figure 1: Left: Illustration of the Yule model using the forward process (“tree growing”). At each step from $n$ to $n+1$, a leaf is chosen uniformly at random and replaced by a cherry. Thus, under the Yule model, the caterpillar tree on four leaves, occurs with probability $P_{Y,4}(T^{cat}_{4})=\frac{2}{3}$, whereas the fully balanced tree of height two occurs with probability $P_{Y,4}(T^{fb}_{2})=\frac{1}{3}$. Note that only considering $n=4$ leaves might lead to the wrong impression that the Yule model favors highly imbalanced trees. In fact, the caterpillar tree is actually getting less and less likely the more leaves are involved (as only the choice of any of the two leaves of the unique cherry in the caterpillar on $n$ leaves can generate a caterpillar on $n+1$ leaves). In fact, as the figure shows, already for $n=5$, the caterpillar is not the most likely tree anymore. Moreover, note that there are 12 phylogenetic trees of caterpillar shape and 3 phylogenetic trees of fully balanced shape. Thus, under the Yule model, the probability to observe a specific phylogenetic tree $\mathcal{T}$ of caterpillar shape is $P_{Y,4}(\mathcal{T})=\frac{2}{3}\cdot\frac{1}{12}=\frac{1}{18}$, and the probability to observe a specific phylogenetic tree $\mathcal{T}$ of fully balanced shape is $P_{Y,4}(\mathcal{T})=\frac{1}{3}\cdot\frac{1}{3}=\frac{1}{9}$.